# 2cm2cm1cm1cm

### v4

@@ 1 Introduction
Quadratic Hamiltonians have been used for decades in theoretical physics, starting with Bogoliubov’s theory of superfluidity [1] and the BCS theory of superconductivity [2, 3, 4, 5] for bosonic and fermionic systems, respectively. They can formally be defined as formal series:
$\mathrm{H}_{0}\doteq\sum_{k,l\in\mathbb{N}}\{\Upsilon_{0}\}_{k,l}a_{k}^{\ast}a_{l}+\{D_{0}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\epsilon\overline{\{D_{0}\}}_{k,l}a_{k}a_{l}+E_{0}\mathbf{1\ ,}$ (1)
@@
By Theorems 2.15 and 2.18 as well as Corollary 2.17, the Hilbert-Schmidt condition on $D_{0}$ can be seen as a Shale-Stinespring condition for generators of strongly continuous unitary groups of Bogoliubov transformations. Recall that the celebrated Shale-Stinespring condition is a sufficient and necessary condition to implement Bogoliubov transformations on the underlying indexing Hilbert space of an abstract CAR $C^{\ast}$–algebras via unitary transformations in the Fock space representation, see [16, Theorems 3.1 and 6.2] or [17, Theorem 7].
@@
Bogoliubov transformations are not only useful to define quadratic Hamiltonians within Bach, Lieb and Solovej’s approach. By definition, they transform creation and annihilation operators in other creation and annihilation operators, leading in theoretical physics to quasi-particle interpretations. Implementable Bogoliubov transformations on Fock spaces are well-known in relatively simple cases to ($\mathrm{N}$–) diagonalize quadratic Hamiltonians $\mathrm{H}_{0}$ of the form (1). The $\mathrm{N}$–diagonalization of $\mathrm{H}_{0}$ refers to the construction of a unitary operator $\mathrm{U}$ acting on the Fock space such that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ commutes with the particle number operator $\mathrm{N}$. It means in this case that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ is a quadratic Hamiltonian of the form (1) with $D_{0}=0$, i.e., the second quantization of a self-adjoint operator acting on the one-particle Hilbert space (see (50)).
@@
Mathematical results on the $\mathrm{N}$–diagonalization of general quadratic operators like (1) have been obtained in [10, 11, 12], but since the sixties no new mathematical result has been performed. Note that Araki presented in 1968 [18] a very general method for the “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians, but it does not solve the issue addressed here. For more details, see Section 2.4.
@@
The second main result of this paper concerns the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians under much more general conditions than before. Compare indeed the previously known result given below by Theorem 2.3, proven in [11, 12], with Theorem 2.4. For instance, observe that the gap condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and the Hilbert-Schmidt condition on the commutator $[\Upsilon_{0},D_{0}]$ in Theorem 2.3 are clearly very strong restrictions as compared to the hypotheses of Theorem 2.4, which can be applied to lower semibounded $\Upsilon_{0}$. In some cases, we are even able to fully characterize the spectrum of fermionic quadratic Hamiltonians. Note that all these results are done by mean of Bogoliubov transformations, implemented in the Fock space, as expected.
@@
The $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians formally defined by (1) may be achieved with presumably optimal conditions in a similar way to what it is efficiently done in [8] for the bosonic case. Here, we follow another path by using the Brockett-Wegner flow [19, 20, 21, 22] like in [6], because this method of proof is interesting in its own right, while leading excellent results. Indeed, such differential equations on spaces of operators are very little developed in Mathematics, being in general very challenging. For instance, although the Brockett-Wegner flow has been often used in theoretical physics [23], there are only two rigorous results on the well-posedness of the Brockett-Wegner flow and only one study of its asymptotics for unbounded operators [6, 21]. By using fermionic quadratic operators we give another example of its well-posedness and asymptotics, by Theorem 3.8, Proposition 3.9 and Corollary 3.12. In fact, the Brockett-Wegner flow [19, 20, 21, 22] is used as a guideline in order to connect the diagonalization of quadratic Hamiltonians to a new, elegant flow on operators acting on the one-particle Hilbert space. In [24] we demonstrate that this non-linear flow presents remarkable ellipticity properties that turn out to be crucial for the study of the infinite-time limit of its solution, which is proven under relatively weak hypotheses on the initial data. It is the elliptic counterpart of an hyperbolic flow used to ($\mathrm{N}$–) diagonalize bosonic quadratic Hamiltonians in [6]. Essential results of [24] for the diagonalization of quadratic Hamiltonians are shortly explained in Section A, while the present paper gives a salient application of this elliptic flow to quantum field theory and quantum statistical mechanics.
@@
To conclude, our main results are the following: (a) Theorem 2.15, and Corollaries 2.14 and 2.17; (b) Theorem 2.4. Observe that (a) and related results belong to the mathematical foundations of fermionic quadratic Hamiltonians, similar to Bruneau and Derezinski’s study [7, 9] on bosonic quadratic Hamiltonians. (b) focuses on their $\mathrm{N}$–diagonalization, similar to the studies [6, 8].
@@
The paper is organized as follows: Section 2 explains fermionic quadratic Hamiltonians, by including our main results while discussing the method of proof via the Brockett-Wegner flow. Most of the technical results as well as a paradigmatic (and historical) example of fermionic quadratic Hamiltonian are gathered in Section 3. In an attempt to be as self-contained and pedagogical as possible, it also contains a review of results that are of general utility in this paper. Section A is an appendix on the elliptical flow studied in [24], which gives the pivotal results used here to diagonalize quadratic Hamiltonians.
@@ 2.1 Preliminary Definitions
The assumption $\mathfrak{h}\doteq L^{2}(\mathcal{M})$ slightly simplifies the setting. More generally, one can take any separable complex Hilbert space $\mathfrak{h}$ endowed with a complex conjugation $\mathcal{C}$, which is an antiunitary involution on $\mathfrak{h}$, i.e., an antilinear mapping $\mathfrak{h}\rightarrow\mathfrak{h}$ such that $\mathcal{C}^{2}=\mathbf{1}$ and
@@ 2.2 Fermionic Quadratic Hamiltonians
Let $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ be any Hilbert-Schmidt operator satisfying $D_{0}=-D_{0}^{\top}$. Take $E_{0}\in\mathbb{R}$ and a self-adjoint operator $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ that is bounded from below on the (separable) Hilbert space $\mathfrak{h}$. Pick some real [^1] Real orthonormal bases are only taken for convenience, in order to slightly simplify our computations, but it is not a necessary condition. Note additionally that such a real basis exists, thanks to Zorn’s lemma. See [26, Lemma 4.201]. orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ in the dense domain $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ of $\Upsilon_{0}$ and, for any $k\in\mathbb{N}$, let $a_{k}\doteq a\left(\varphi_{k}\right)$ be the corresponding annihilation operator acting on the fermionic Fock space
@@
the scalar product and the norm of which are respectively denoted by $\left\langle\cdot,\cdot\right\rangle_{\mathcal{F}}$ and $\left\|\cdot\right\|_{\mathcal{F}}$. Note that this scalar product is the sum over $n\in\mathbb{N}$ of each canonical scalar product on the sector $\wedge^{n}\mathfrak{h}$. Here, $\wedge^{0}\mathfrak{h}\doteq\mathbb{C}$, while, for $n\in\mathbb{N}$, $\wedge^{n}\mathfrak{h}$ is the subspace of totally antisymmetric $n$–particle wave functions in $\mathfrak{h}^{\otimes n}$, the $n$–fold tensor product of $\mathfrak{h}$.
For any fixed $E_{0}\in\mathbb{R}$, the fermionic quadratic operator is defined through the operators $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ by
@@
with $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the subspace of totally antisymmetric $n$–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the $n$–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$. We use again the notation $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ for its (uniquely defined) self-adjoint extension. Note additionally that $\mathrm{H}_{0}$ is independent of the (possibily real) orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ chosen, as explained after Proposition 3.4. For a general form of the quadratic Hamiltonian $\mathrm{H}_{0}$ written within an arbitrary orthonormal basis in $\mathcal{D}\left(\Upsilon_{0}\right)$, see Equation (56). Note also that a continuous integral $\int\mathrm{d}k\ \mathrm{d}\ell$ could also replace the discrete sum $\sum_{k,\ell}$ in (4), with appropriate modifications, since all results also hold in the continuous case.
@@
Hamiltonians like (4) appear very often in theoretical physics and can for instance encode in some sense interaction effects. In fact, interparticle interactions usually play a pivotal role in physical phenomena, but they make the physical system extremely difficult to study in a mathematically rigorous way and effective models coming from different approximations and ansäzte are usually used. It is exactly what happens for the celebrated BCS theory proposed in the late 1950s to explain conventional superconductivity, which leads to the explicit diagonalization of a (very simple) quadratic operators like (4). See Section 3.7 for more details. Another example is very recently given in the context of the (generalized) Hartree-Fock theory [15], which has a strong affinity with (fermionic) mean-field models and effective quadratic Hamiltonians, as first discussed in [26, Sections 5.7 and 6.10] (with a research paper in preparation).
@@
In the more general cases where $D_{0}\neq 0$, one can usually find a unitary transformation $\mathrm{U}$ such that $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ is $\mathrm{N}$–diagonal. This was first shown for fermions in the particular case of the BCS theory (Section 3.7) by implementing [^2] Bogoliubov transformations on CAR $C^{\ast}$-algebras are not necessarily implementable on the Fock space, i.e., they do not necessarily well-define a unitary transformation on the Fock space. This is discussed in detail in Section 2.4, see in particular the Shale-Stinespring theorem (Theorem 2.8). on the Fock space an algebraic transformation, named the Bogoliubov (unitary) transformation, which defines a $\ast$-isomorphism of CAR $C^{\ast}$-algebras.
@@ 2.3 Diagonalization of Quadratic Hamiltonians
Let $E_{0}\in\mathbb{R}$. Take $D_{0}=\bar{D}_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\top}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with strictly positive $\alpha\in\mathbb{R}^{+}$. Assume additionally that the commutator $[D_{0},\Upsilon_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ and
@@
Note that Araki presented in 1968 [18] a very general method for the “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians. The results are done in great generality by using an algebraic approach, but they do not solve the issue addressed here. For more details, we recommend Section 2.4, where Araki’s work is described in detail.
In fact, after 1967, to our knowledge, there is surprisingly no new general results on the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians described here. The same situation appears for the bosonic case for which no new results on such Hamiltonians were performed since the 1960ies [10, 11, 12], until the year 2007 with the paper [7]. Then, driven by the mathematical justification of Bogoliubov’s theory or the Brockett-Wegner flow, new results on such models have been developed in recent years, see [6, 8, 9], with for instance the efficient proof on the $\mathrm{N}$–diagonalization of bosonic quadratic Hamiltonians under presumably optimal conditions [8].
By contrast, there is no recent activity on the (equally important) fermionic case and we address the issue of the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians by means of an elliptic operator-valued flow (Section A). It leads to a continuous family of Bogoliubov transformations and gives new properties on the $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$, allowing in some case to compute it explicitly. This is explained in Section 2.5 while here we only sum up the final result on the $\mathrm{N}$–diagonalization of quadratic Hamiltonians:
@@
for some $\0\mu\in\mathbb{R}\backslash\{0\}$ and $\varepsilon\in\mathbb{R}^{+}$. Then, there is a unitary transformation $\mathrm{U}$ such that
@@
To get a stronger exponential decay of the Hilbert-Schmidt norm $\left\|D_{t}\right\|_{2}$ and a more accurate estimate of the spectral gap of $\Upsilon_{\infty}$, one shall take the maximum value of $\mu\in\mathbb{R}$ such that (13) holds true (provided such a $\mu$ exists).
By Proposition 3.6 and Corollary 3.11, the unitary transformation $\mathrm{U}$ implements a Bogoliubov transformation. The $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$ can be additionally computed in some cases. For instance, using the constant of motion (192) and Theorem A.2 with $D_{0}^{\top}=\pm D_{0}$, we have
@@
at least if $\Upsilon_{0}$ is bounded. In the case that $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$ on the domain $\mathcal{D}(\Upsilon_{0}^{\top})$ together with $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\mu\in\mathbb{R}$ and $\varepsilon\in\mathbb{R}^{+}$ and
@@
for some strictly positive constant $\mathrm{C}\in\mathbb{R}^{+}$. Then, [24, Equation (11)] yield
@@
Note that the fact that the trace in the above equation is finite is a non-trivial [^3] For a direct proof under the condition $\Upsilon_{0}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$, see [11, Lemma 8.1]. consequence of the study. In other words, $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ can be fully characterized in this “commutative” case. This is for instance used in Section 3.7 to recover the exact $\mathrm{N}$–diagonal form (187) of the BCS Hamiltonian.
Note that Theorem 2.3 does not make explicit neither the $\mathrm{N}$–diagonal form of fermionic quadratic Hamiltonians nor the Bogoliubov transformation. In addition, observe that (13) looks similar to Condition (11) of Theorem 2.3, but the assumptions $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and $[\Upsilon_{0},D_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.3 are clearly strong restrictions as compared to the hypotheses of Theorem 2.4. See for instance Remark 2.5. In fact, the asssumptions of Berezin’s theorem (Theorem 2.3) are so strong that they may not even be satisfied by the elementary quadratic Hamiltonians of the BCS theory, as discussed in Section 3.7. Indeed, for these Hamiltonians, the condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ may not necessarily hold for a strictly positive $\alpha\in\mathbb{R}^{+}$. Meanwhile, the assumptions of Theorem 2.4 are much more natural and general than those of Theorem 2.3 and they include important cases with negative spectra for $\Upsilon_{0}$.
To conclude, note that the conditions of Theorems 2.4 are only sufficient. As already discussed in [24] we expect the limiting case to be when

### v6

# Quadratic Hamiltonians in Fermionic Fock Spaces

@@ 1 Introduction
Quadratic Hamiltonians have been used for decades in theoretical physics, starting with Bogoliubov’s theory of superfluidity [1] and the BCS theory of superconductivity [2, 3, 4, 5] for bosonic and fermionic systems, respectively. They can formally be defined (up to a certain constant) as formal series:
$\mathrm{H}_{0}\doteq\sum_{k,l\in\mathbb{N}}\{\Upsilon_{0}\}_{k,l}a_{k}^{\ast}a_{l}+\{D_{0}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\epsilon\overline{\{D_{0}\}}_{k,l}a_{k}a_{l}\mathbf{\ ,}$ (1)
@@
The Shale-Stinespring condition on Bogoliubov transformations does not necessarily imply a Hilbert-Schmidt condition on the off-diagonal elements of its generator, but, by Theorems 2.15 and 2.18 as well as Corollary 2.17, the Hilbert-Schmidt condition on $D_{0}$ can be seen as Shale-Stinespring-like condition for quadratic operators. Recall that the celebrated Shale-Stinespring condition is a sufficient and necessary condition to implement Bogoliubov transformations on the underlying indexing Hilbert space of an abstract CAR $C^{\ast}$–algebras via unitary transformations in the Fock space representation, see [16, Theorems 3.1 and 6.2] or [17, Theorem 7].
@@
Bogoliubov transformations are not only useful to define quadratic Hamiltonians within Bach, Lieb and Solovej’s approach. By definition, they transform creation and annihilation operators in other creation and annihilation operators, leading in theoretical physics to quasi-particle interpretations. Implementable Bogoliubov transformations on Fock spaces are well-known in relatively simple cases to ($\mathrm{N}$–) diagonalize quadratic Hamiltonians $\mathrm{H}_{0}$ of the form (1). The $\mathrm{N}$–diagonalization of $\mathrm{H}_{0}$ refers to the construction of a unitary operator $\mathrm{U}$ acting on the Fock space such that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ commutes with the particle number operator $\mathrm{N}$. It means in this case that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ is a quadratic Hamiltonian of the form (1), up to a certain constant, with $D_{0}=0$, i.e., the second quantization of a self-adjoint operator acting on the one-particle Hilbert space (see (47)).
@@
Mathematical results on the $\mathrm{N}$–diagonalization of general quadratic operators like (1) have been obtained in [10, 11, 12], but since the sixties no new mathematical result has been performed. Note that Araki presented in 1968 [18] a very general method for the $\mathrm{N}$–diagonalization” of bilinear Hamiltonians, but it does not solve the issue addressed here. For more details, see Section 2.4.
@@
The second main result of this paper concerns the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians under much more general conditions than before. Compare indeed the previously known result given below by Theorem 2.3, proven in [11, 12], with Theorem 2.4. For instance, observe that the gap condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and the Hilbert-Schmidt condition on the commutator $\Upsilon_{0},D_{0}]$ in Theorem 2.3 are clearly very strong restrictions as compared to the hypotheses of Theorem 2.4, which can be applied to lower semibounded $\Upsilon_{0}$. In some cases, we are even able to fully characterize the spectrum of fermionic quadratic Hamiltonians. Note that all these results are done by means of Bogoliubov transformations, implemented in the Fock space, as expected.
@@
While aware of the method used in [8] for the bosonic case, we do no try to use it to handle the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians formally defined by (1). Here, we follow another path by using the Brockett-Wegner flow [19, 20, 21, 22] like in [6], because [8] does not necessarily give explicit expressions for the $\mathrm{N}$-diagonal form of the Hamiltonian and, in any case, the Brockett-Wegner method of proof is interesting in its own right, while leading excellent results like explicit expressions in special cases. Indeed, such differential equations on spaces of operators are very little developed in Mathematics, being in general very challenging. For instance, although the Brockett-Wegner flow has been often used in theoretical physics [23], there are only two rigorous results on the well-posedness of the Brockett-Wegner flow and only one study of its asymptotics for unbounded operators [6, 21]. By using fermionic quadratic operators we give another example of its well-posedness and asymptotics, by Theorem 3.8, Proposition 3.9 and Corollary 3.12. In fact, the Brockett-Wegner flow [19, 20, 21, 22] is used as a guideline in order to connect the diagonalization of quadratic Hamiltonians to a new, elegant flow on operators acting on the one-particle Hilbert space. In [24] we demonstrate that this non-linear flow presents remarkable ellipticity properties that turn out to be crucial for the study of the infinite-time limit of its solution, which is proven under relatively weak hypotheses on the initial data. It is the elliptic counterpart of an hyperbolic flow used to ($\mathrm{N}$–) diagonalize bosonic quadratic Hamiltonians in [6]. Essential results of [24] for the diagonalization of quadratic Hamiltonians are shortly explained in Section A.1, while the present paper gives a salient application of this elliptic flow to quantum field theory and quantum statistical mechanics.
@@
To conclude, our main results are the following: (a) Theorem 2.15, and Corollaries 2.14 and 2.17; (b) Theorem 2.4. Observe that (a) and related results belong to the mathematical foundations of fermionic quadratic Hamiltonians, similar to Bruneau and Derezinski’s study [7, 9] on bosonic quadratic Hamiltonians. (b) focuses on their $\mathrm{N}$–diagonalization, as is studied in [6, 8] in the bosonic case. Comparatively to [8] done in the bosonic case, we obtain here some expressions for the diagonal form as integrals of elements of the elliptic flow, and in some cases even very explicit expressions (see (16) and (17)).
@@
The paper is organized as follows: Section 2 explains fermionic quadratic Hamiltonians, by including our main results while discussing the method of proof via the Brockett-Wegner flow. Most of the technical results are gathered in Sections 3 and 4. Inside the Appendix A, we briefly expose in Section A.1 the pivotal results used on the elliptical flow studied in [24], while in Section A.2 we give a paradigmatic (and historical) example of fermionic quadratic Hamiltonian.
@@ 2.1 Preliminary Definitions
The hypothesis $\mathfrak{h}\doteq L^{2}(\mathcal{M})$ is used solely for the sake of simplicity and, to be specific, is in fact only used in Equation (163). So, one can simply take any separable complex Hilbert space $\mathfrak{h}$ endowed with a complex conjugation $\mathcal{C}$, which is an antiunitary involution on $\mathfrak{h}$, i.e., an antilinear mapping $\mathfrak{h}\rightarrow\mathfrak{h}$ such that $\mathcal{C}^{2}=\mathbf{1}$ and
@@ 2.2 Fermionic Quadratic Hamiltonians
Let $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ be any Hilbert-Schmidt operator satisfying $D_{0}=-D_{0}^{\top}$. Take $E_{0}\in\mathbb{R}$ and a self-adjoint operator $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ that is bounded from below on the (separable) Hilbert space $\mathfrak{h}$. Pick some real  [^1] In the general setting, a real orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ means that $\varphi_{k}=\mathcal{C}\varphi_{k}$ for all $k$. If $\mathfrak{h}$ can be decomposed as $\mathfrak{h}=\mathfrak{X}\oplus\mathfrak{X}^{\ast}$ for some Hilbert space $\mathfrak{X}$, then such a real basis always exists, thanks to Zorn’s lemma. See [26, Lemma 4.200]. In any case, real orthonormal bases are not required, but we assume this situation to slightly lighten the computations. orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ in the dense domain $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ of $\Upsilon_{0}$ and, for any $k\in\mathbb{N}$, let $a_{k}\doteq a\left(\varphi_{k}\right)$ be the corresponding annihilation operator acting on the fermionic Fock space
@@
the scalar product and the norm of which are respectively denoted by $\left\langle\cdot,\cdot\right\rangle_{\mathcal{F}}$ and $\left\|\cdot\right\|_{\mathcal{F}}$. Note that this scalar product is the sum over $n\in\mathbb{N}$ of each canonical scalar product on the sector $\wedge^{n}\mathfrak{h}$. Here, $\wedge^{0}\mathfrak{h}\doteq\mathbb{C}$, while, for $n\in\mathbb{N}$, $\wedge^{n}\mathfrak{h}$ is the subspace of totally antisymmetric nn–particle wave functions in $\mathfrak{h}^{\otimes n}$, the nn–fold tensor product of $\mathfrak{h}$.
For any fixed $E_{0}\in\mathbb{R}$, the fermionic quadratic operator  [^2] One could simply take $E_{0}=0$, but since this constant will become time-dependent and non-zero later on, we use it in the definition. is defined through the operators $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ by
@@
with $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the subspace of totally antisymmetric nn–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the nn–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$. We use again the notation $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ for its (uniquely defined) self-adjoint extension. Note additionally that $\mathrm{H}_{0}$ is independent of the (possibily real) orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ chosen, as explained after Proposition 3.4. For a general form of the quadratic Hamiltonian $\mathrm{H}_{0}$ written within an arbitrary orthonormal basis in $\mathcal{D}\left(\Upsilon_{0}\right)$, see Equation (52).
@@
Hamiltonians like (4) appear very often in theoretical physics and can for instance encode in some sense interaction effects. In fact, interparticle interactions usually play a pivotal role in physical phenomena, but they make the physical system extremely difficult to study in a mathematically rigorous way and effective models coming from different approximations and ansäzte are usually used. It is exactly what happens for the celebrated BCS theory proposed in the late 1950s to explain conventional superconductivity, which leads to the explicit diagonalization of a (very simple) quadratic operators like (4). See Section A.2 for more details. Another example is given in the context of the (generalized) Hartree-Fock theory [15], which has a strong affinity with (fermionic) mean-field models and effective quadratic Hamiltonians, as first discussed in [26, Sections 5.7 and 6.10] (with a research paper in preparation).
@@
In the more general cases where $D_{0}\neq 0$, one can usually find a unitary transformation $\mathrm{U}$ such that $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ is $\mathrm{N}$–diagonal. This was first shown for fermions in the particular case of the BCS theory (Section A.2) by implementing [^3] Bogoliubov transformations on CAR $C^{\ast}$-algebras are not necessarily implementable on the Fock space, i.e., they do not necessarily well-define a unitary transformation on the Fock space. This is discussed in detail in Section 2.4, see in particular the Shale-Stinespring theorem (Theorem 2.8). on the Fock space an algebraic transformation, named the Bogoliubov (unitary) transformation, which defines a $\ast$-isomorphism of CAR $C^{\ast}$-algebras.
@@ 2.3 Diagonalization of Quadratic Hamiltonians
Let $E_{0}\in\mathbb{R}$. Take $D_{0}=\bar{D}_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\top}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with strictly positive $\alpha\in\mathbb{R}^{+}$. Assume additionally that the commutator $D_{0},\Upsilon_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ and
@@
Note that Araki presented in 1968 [18] a very general method for the $\mathrm{N}$–diagonalization” of bilinear Hamiltonians. The results are done in great generality by using an algebraic approach, but they do not solve the issue addressed here. For more details, we recommend Section 2.4, where Araki’s work is described in detail.
In fact, after 1967, to our knowledge, there is surprisingly no new general results on the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians described here. The same situation appears for the bosonic case for which no new results on such Hamiltonians were performed since the 1960ies [10, 11, 12], until the year 2007 with the paper [7]. Then, driven by the mathematical justification of Bogoliubov’s theory or the Brockett-Wegner flow, new results on such models have been developed in recent years, see [6, 8, 9], as on the $\mathrm{N}$–diagonalization of bosonic quadratic Hamiltonians [6, 8].
By contrast, there is no recent activity on the (equally important) fermionic case and we address the issue of the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians by means of an elliptic operator-valued flow (Section A.1). It leads to a continuous family of Bogoliubov transformations and gives new properties on the $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$, allowing in some case to compute it explicitly. This is explained in Section 2.5 while here we only sum up the final result on the $\mathrm{N}$–diagonalization of quadratic Hamiltonians:
@@
[…]
@@
In practice, one shall take the maximum value of $\mu\in\mathbb{R}$ such that (13) holds true (provided such a $\mu$ exists) in order to obtain a more accurate estimate of the spectral gap of $\Upsilon_{\infty}$.
By Proposition 3.6 and Corollary 3.11, the unitary transformation $\mathrm{U}$ implements a Bogoliubov transformation. The $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$ can be additionally computed in some cases. For instance, using the constant of motion given by [24, Equation (23)] and Theorem A.2 with $D_{0}^{\top}=\pm D_{0}$, we have
@@
at least if $\Upsilon_{0}$ is bounded. If $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$ on the domain $\mathcal{D}(\Upsilon_{0}^{\top})$, $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\mu\in\mathbb{R}$ and $\varepsilon\in\mathbb{R}^{+}$ and
@@
for some strictly positive constant $\mathrm{C}\in\mathbb{R}^{+}$, then [24, Equation (11)] yield
@@
Note that the fact that the trace in the above equation is finite is a non-trivial [^4] For a direct proof under the condition $\Upsilon_{0}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$, see [11, Lemma 8.1]. consequence of the study. In other words, $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ can be fully characterized in this “commutative” case. This is for instance used in Section A.2 to recover the exact $\mathrm{N}$–diagonal form (184) of the BCS Hamiltonian.
Note that Theorem 2.3 does not make explicit neither the $\mathrm{N}$–diagonal form of fermionic quadratic Hamiltonians nor the Bogoliubov transformation. In addition, observe that (13) looks similar to Condition (11) of Theorem 2.3, but the assumptions $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and $\Upsilon_{0},D_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.3 are clearly strong restrictions as compared to the hypotheses of Theorem 2.4. See for instance Remark 2.5. In fact, the asssumptions of Berezin’s theorem (Theorem 2.3) are so strong that they may not even be satisfied by the elementary quadratic Hamiltonians of the BCS theory, as discussed in Section A.2. Indeed, for these Hamiltonians, the condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ may not necessarily hold for a strictly positive $\alpha\in\mathbb{R}^{+}$. Meanwhile, the assumptions of Theorem 2.4 are much more natural and general than those of Theorem 2.3 and they include important cases with negative spectra for $\Upsilon_{0}$.
To conclude, note that the conditions of Theorem 2.4 are not necessary (in principle). As already discussed in [24] we expect the limiting case to be when [^5] X>0X>0 means $X\geq 0$, with 0 not being in the spectrum of $X$.